Let f(x) = (x^2−1)/(x^4+1)

(a) At which points does the graph of the f(x) have a horizontal tangent line?

(b) Draw the graph of f(x) on MATLAB(or octave online) and identify the points for horizontal tangents on the graph.


1
Expert's answer
2021-10-18T15:22:17-0400

a)

horizontal tangent line is at points of extremum:

"f'(x)=\frac{2x(x^4+1)-4x^3(x^2-1)}{(x^4+1)^2}=0"

"2x^5+2x-4x^5+4x^3=0"

"x(2-2x^4+4x^2)=0"

"x_1=0"

let "k=x^2" , then:

"k^2-2k+1=0"

"k=1"

"x_2=1,x_3=-1"


b)

x=-100:100;
y=(x.^2-1)./(x.^4+1);
plot(x,y)